# Amnesiac Flooding: Easy to break, hard to escape

### v1

@@ Acknowledgements.
If there does not exist $k\leq 2|E|$ such that $A^{k}(S)=\emptyset$ then there must be a message path of length $12|E|+$ 1 for some message in $S$. By the pigeon hole principle this message path must visit the same edge twice travelling in the same direction. Thus, this requires a recurrent message. Therefore, $S$ cannot be balanced. ∎
Appendix B Proof of Theorem 4.4
@@
    For $u,v,w\in[\kappa]$ if $f(u,\{v\},\{v\})=\{v\}$ and $f(w,\{v\},\{v\})=\{v\}$ then either $f(v,\{u,w\},\{u\})=\{u,w\}$, $f(v,\{u,w\},\{w\})=\{u,w\}$ or both. Additionally, $f(v,\{u,w\},\{u,w\})=\emptyset$. Otherwise the protocol would not terminate on the three node path with labels $uvw$.
@@
We begin by considering $u,v,w,x,y,z\in[\kappa]$ such that $u>_{*}v<_{*}w$ and $x>_{*}y<_{*}z$. By Observation B.1 (4), wlog. we can assume that $f(v,\{u,w\},\{u\})=\{u,w\}$ and that $f(y,\{x,z\},\{z\})=\{x,z\}$. Now consider the path on 6 nodes labelled $uvwxyz$, as well as two paths on three nodes labeled $uvw$ and $xyz$. Since $u,v,y,z$ have the same neighbourhood on $P_{6}$ as on their corresponding $P_{3}$ they must be obeying the same policy. Thus, only $w$ and $x$ have a different policy. However, by exhaustive search it can be shown that there exists no protocol that will terminate from every node of the $P_{6}$, as well as each connected sub-graph. Thus there cannot exist such $u,v,w,x,y,z$.
@@
If there exists $S\subseteq[\kappa]$ such that $|S|>6$ where $\forall u,v\in S$ $u\not>_{*}v$, then there exists $T\subseteq S$ with $|T|=|S|-3$ such that $\forall u,v,w\in T$, $f(v,\{u,w\},\{u\})=\{w\}$.
@@
There are several cases to consider. If there are no triples of the form $u,v,w\in S$ such that $f(v,\{u,w\},\{u\})=\{u,w\}$ then the claim is trivial. Now assume that all such triples include one of the three identifiers $x,y,z\in S$. In this case, we take the remaining $|S|-3$ to be $T$ and the claim holds. Finally, assume that there are two disjoint triples $a,b,c$ and $g,h,i$ with this property. Without loss of generality we can assume that $f(b,\{a,c\},\{a\})=\{a,c\}$ and $f(h,\{g,i\},\{g\})=\{g,i\}$. Taking the path on $6$ nodes labelled $abcghi$, we have that either $b,c,g$ or $c,g,h$ must also be such a triple, or the protocol will not terminate on our path. However, should $bcg$ or $cgh$ be such triples, we can shorten the path by deleting vertices from each end to obtain a labelling that once again does not terminate. More concretely, while ever there are only two such triples on the path that are “facing each other” they will bounce the message back and forth between each other. If there are more than two triples we just remove vertices until only two remain. Thus, as $P$ terminates for all graphs and labellings there of, we have that all such triples share one of three identifiers, and so we can take $T$ to be $S\setminus\{x,y,z\}$. ∎
@@
With this established we can extend the argument to all labels. While technically the $T$ referred to in the previous lemma is not necessarily $[|T|]$, all labels within $[\kappa]$ can be reordered arbitrarily as they are available for all graphs. For this reason and ease of presentation we will simply take $[T]=[|T|]$.
@@
We now need to show that $f(\alpha,\{v,\beta\},\{v\})=\{\beta\}$ as well. Consider a path on $m+1$ nodes labelled $wxvyzc...d$, where the section $c...d$ contains all remaining labels from $[m]$. Since, only $v,w,y$ could have policies different to those of Amnesiac Flooding in this setting, a message can only be sent from outside the subsection $v,w,y$ once, and so we can treat this structure like $P_{3}$. Similar to the path on three nodes, without loss of generality, there are essentially 12 scenarios to consider, which produce behaviours falling into 5 general cases expressed in the table in figure 7.
@@
It remains to construct a communication graph and labelling such that for each case $1$-$4$ termination fails.
For case $1$, there is exactly one of $x,v,y$ which implements a policy different to Amnesiac Flooding, we denote it $\gamma$. Furthermore, without loss of generality there exists at least one of its neighbours $\beta$ such that upon receiving a message from $\beta$ it sends messages to both of its neighbours. We take the path $wxvyz$, which contains either the segment $\beta\gamma$ or $\gamma\beta$. In the former case, we attach $a$ and $b$ to $w$ to form a $C_{3}$ on $abw$ and the path $c...d$ to $z$, forming an extended paw graph. In the latter case, we instead attach $a$ and $b$ to $z$ and the path on $c...d$ to $w$ to again form an extended paw. Now consider broadcast initiated at the solitary leaf node. The message will reach $\gamma$ and then bounce back and forth between $\gamma$ and the $C_{3}$ for ever.
@@
Figure 8. A communication graph for Case 2 where the path $c...d$ contains all remaining identifiers from $[m]$ and a non-terminating message sequence initiated at node $z$. The message sequence is given to the final unique configuration.
@@
Figure 9. Non-terminating messages sequences for cases $3$ and $4$ in lemma B.7. Both behave identically for the first 9 iterations, however case $3$ takes two further iterations to repeat a step (indicated by dashed lines). Note that path $c...d$ has been omitted here but would be attached to $z$.
Thus as we have obtained non-termination for all cases other than case 00, we must have for all $\alpha,\beta,\gamma\in[m+1]$ that $f(\alpha,\{\beta,\gamma\},\{\beta\})=\{\gamma\}$. Furthermore, we must have that $b(\alpha,\{\beta,\gamma\})=\{\beta,\gamma\}$ as otherwise the protocol would not terminate on a cycle. This also implies that $f(\alpha,\{\beta,\gamma\},\{\beta,\gamma\})=\emptyset$ to ensure termination on odd cycles when broadcast is initiated at $\alpha$. Finally, we must have that $f(\alpha,\{\beta\},\{\beta\})=\emptyset$ as $\alpha$ could be on the end of an extended paw. Thus, we have that $P$ is AF on $[m+1]$ up to degree $2$.
Fortunately, we can reuse much of the previous argument for obtaining degree $3$ behaviour. Excluding the diamond graph and the $K_{3,2}$ graph, we can use all of the constructions in Lemma B.5 by simply adding a path $c...d$ to a node of degree at most $2$ which is not adjacent to $v$. However, we can slightly alter the diamond and $K_{3,2}$ graph to permit the addition of such path to a node of degree at most $2$ which is not adjacent to $v$ no matter where $v$ appears in the labelling. These are given explicitly in figure Fig. 10. Note that while, these do require $8$ nodes to construct, we can use the original constructions from Lemma B.5 until we have more than $\kappa$ nodes.
@@
Figure 11. Two graphs used in the proof of theorem 4.4 to determine identifiers’ response to receiving only a single message. In both cases $c...d$ is a path containing all unused identifiers from $[m]$ where $m$ is the highest id used in labelling the body of the graph. Left: A tree that forbids sending to too small a subset of neighbours. Right: A graph that forbids sending a message to all neighbours.
@@
Figure 12. A graph used in the proof of theorem 4.4. The graph consists of a star centred at $u$ with its leaves partitioned into two sets of size $r$ and $k-r$. The leaves in the set of size $r$ are connected by a binary tree of depth $\lceil\log_{2}{r}\rceil$ with root $w^{*}$ to a cycle, which in turn is connected to a path. Note that the path $c...d$ contains all identifiers from $[m]$ not used in the labelling of the rest of the graph, where $m$ is the largest $id$ present.
@@
Figure 13. A graph used in the proof of Theorem 4.4. The graph consists of a star centred at $u$ with its leaves partitioned into three sets of size $r$, $q-r+1$ and $k-q-1$. The leaves in the first two sets are then each joined to a single node labelled by $w*$ and $v*$ respectively by binary trees of depth $\lceil\log_{2}{q}\rceil$. The single nodes are connected to a node labelled $a$ which is the start of a path $ac...d$ that contains all identifiers in $[m]$ not used in the labelling of the body of the graph where $m$ is the largest id used in that labelling.
@@ Appendix C Uniqueness: Existence of Protocols in the relaxed cases
In this appendix, we present protocols for each of the relaxations discussed in Section 5
@@ C.2. Protocol with non-determinism
Random Flooding is the protocol defined by the following rule: on each round, if a node has a random bit equal to 00 it implements Amnesiac Flooding. Otherwise upon receiving a message it sends to all neighbours.
@@
Consider the random variable $X(t)$ which for $t\in\mathbb{N}$ takes the value $1$ if on rounds $t,t+1,...,t+D$ all nodes receive $1$ as a random bit, but at round $t+D+1$ all nodes receive 00, and 00 otherwise.
@@ C.3. Protocols that can read messages/send multiple messages
1-Bit Flooding is the protocol defined according to the following rule: When broadcast begins, the initiator picks 00 for the shared bit if it is a leaf and $1$ otherwise. Upon receiving a message, if the shared bit is a $1$ the node implements parrot flooding, and implements Amnesiac Flooding otherwise.
@@
If we have readable header information, then the single bit is placed in the header. On the other hand, if the nodes are able to send multiple messages per edge per round, then the initiator may choose to send $2$ messages in the case where it wishes to encode a $1$ and a single one otherwise. In this case, if nodes always forward the same number of messages then the single bit is propagated and maintained.
@@
Let $S$ be a configuration of messages on $G=(V,E)$. Since messages are only reintroduced at leaves rather than on cycles or FECs, $P(S)$ is balanced on $G$ if and only if $S$ is balanced on $G$. Formally, this follows from Theorem.
However, if parrot flooding is initiated at the terminal vertex of a path this configuration is balanced but will not terminate. Therefore, we require a stronger condition than balance to capture termination under parrot flooding. Consider a message path of parrot flooding (the definition of which is obtained by simply replacing the operator $A$ with $P$ in definition A.3). We immediately inherit Lemmas A.4, A.5, A.6, A.6 and A.8 as none of them depend on the behaviour of leaves. Thus any recurrent message $(u,v)\in S$ must have a message path $uv...wlw...uv$ where $l$ is a leaf node or it would be captured by the argument used in the proof of Theorem 6.1. Therefore, in fact there exists a configuration $\hat{S}=P^{k}(S)$ for some $k\in\mathbb{N}$ such that $(w,l)\in\hat{S}$ where $(w,l)$ is recurrent on $\hat{S}$. The rest of the argument is similar to that used in the proof of Theorem 6.1, and we will construct a suitable notion of balance. First, however, we must define the additional structure that produces non-termination.
@@
We have already established that balance is preserved by parrot flooding, so it remains to show that $P$-Balance is conserved on all leaf paths. Let $Q$ be a leaf path starting at $u$ and ending at $v$ on $G=(V,E)$ and let $S$ be a configuration on $G$. Then $P_{G}(S)_{Q_{p}}=P_{Q_{p},I}(S)$ for some $I\subset V$ such that $u,v\not\in I$, i.e. a step of parrot flooding on $G$ is equivalent to a step of parrot flooding on the leaf path representation with some possible external node activation. After a step of parrot flooding, two messages on $Q_{p}$ will either have the distance between their heads: stay the same, increase by two or decrease by two. Upon externally activating a non-leaf node two messages are created, with their heads an even distance apart on $Q_{p}$. Symmetrically, a collision between two messages can only be between two messages with an even distance (explicitly 00) between their heads and removes both. Thus, $P$-Balance is maintained. ∎
@@ Appendix D Proof of the Fault Sensitivity results
In this appendix, we prove the fault sensitivity results of section Section 6, using Theorem 6.1 as our main technical tool. We begin with the single message failure case, restated below.
Theorem 0 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate. if either:
  • $uv$ is not a bridge
  • $uv$ lies on a FEC subgraph
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, $uv$ is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.

### v2

@@ 6. Conclusions and Future Work
If there does not exist $k\leq 2|E|$ such that $A^{k}(S)=\emptyset$ then there must be a message path of length $2|E|+1$ for some message in $S$. By the pigeon hole principle this message path must visit the same edge twice travelling in the same direction. Thus, this requires a recurrent message. Therefore, $S$ cannot be balanced. ∎
Appendix B Proof of Uniqueness (Theorem 4.3)
@@
    For $u,v,w\in[\kappa]$ if $f(u,\{v\},\{v\})=\{v\}$ and $f(w,\{v\},\{v\})=\{v\}$ then either $f(v,\{u,w\},\{u\})=\{u,w\}$, $f(v,\{u,w\},\{w\})=\{u,w\}$ or both. Additionally, $f(v,\{u,w\},\{u,w\})=\emptyset$. Otherwise the protocol would not terminate on the three node path with labels uvwuvw.
@@
We begin by considering $u,v,w,x,y,z\in[\kappa]$ such that $u>_{*}v<_{*}w$ and $x>_{*}y<_{*}z$. By Observation B.1 (4), wlog. we can assume that $f(v,\{u,w\},\{u\})=\{u,w\}$ and that $f(y,\{x,z\},\{z\})=\{x,z\}$. Now consider the path on 6 nodes labelled uvwxyzuvwxyz, as well as two paths on three nodes labeled uvwuvw and xyzxyz. Since $u,v,y,z$ have the same neighbourhood on $P_{6}$ as on their corresponding $P_{3}$ they must be obeying the same policy. Thus, only $w$ and $x$ have a different policy. However, by exhaustive search it can be shown that there exists no protocol that will terminate from every node of the $P_{6}$, as well as each connected sub-graph. Thus there cannot exist such $u,v,w,x,y,z$.
@@
If there exists $S\subseteq[\kappa]$ such that $|S|>6$ where $\forall u,v\in S$ $u\not>_{*}v$, then there exists $T\subseteq S$ with |T|=|S|3|T|=|S|-3 such that $\forall u,v,w\in T$, $f(v,\{u,w\},\{u\})=\{w\}$.
@@
There are several cases to consider. If there are no triples of the form $u,v,w\in S$ such that $f(v,\{u,w\},\{u\})=\{u,w\}$ then the claim is trivial. Now assume that all such triples include one of the three identifiers $x,y,z\in S$. In this case, we take the remaining |S|3|S|-3 to be $T$ and the claim holds. Finally, assume that there are two disjoint triples $a,b,c$ and $g,h,i$ with this property. Without loss of generality we can assume that $f(b,\{a,c\},\{a\})=\{a,c\}$ and $f(h,\{g,i\},\{g\})=\{g,i\}$. Taking the path on $6$ nodes labelled abcghiabcghi, we have that either $b,c,g$ or $c,g,h$ must also be such a triple, or the protocol will not terminate on our path. However, should bcgbcg or cghcgh be such triples, we can shorten the path by deleting vertices from each end to obtain a labelling that once again does not terminate. More concretely, while ever there are only two such triples on the path that are “facing each other” they will bounce the message back and forth between each other. If there are more than two triples we just remove vertices until only two remain. Thus, as $P$ terminates for all graphs and labellings there of, we have that all such triples share one of three identifiers, and so we can take $T$ to be $S\setminus\{x,y,z\}$. ∎
@@
With this established we can extend the argument to all labels. While technically the $T$ referred to in the previous lemma is not necessarily $[|T|]$, all labels within $\kappa]$ can be reordered arbitrarily as they are available for all graphs. For this reason and ease of presentation we will simply take $[T]=[|T|]$.
@@
We now need to show that $f(\alpha,\{v,\beta\},\{v\})=\{\beta\}$ as well. Consider a path on $m+1$ nodes labelled wxvyzcdwxvyzc...d, where the section cdc...d contains all remaining labels from $[m]$. Since, only $v,w,y$ could have policies different to those of Amnesiac Flooding in this setting, a message can only be sent from outside the subsection $v,w,y$ once, and so we can treat this structure like $P_{3}$. Similar to the path on three nodes, without loss of generality, there are essentially 12 scenarios to consider, which produce behaviours falling into 5 general cases expressed in the table in figure 7.
@@
It remains to construct a communication graph and labelling such that for each case 11-44 termination fails.
For case $1$, there is exactly one of $x,v,y$ which implements a policy different to Amnesiac Flooding, we denote it $\gamma$. Furthermore, without loss of generality there exists at least one of its neighbours $\beta$ such that upon receiving a message from $\beta$ it sends messages to both of its neighbours. We take the path wxvyzwxvyz, which contains either the segment $\beta\gamma$ or $\gamma\beta$. In the former case, we attach $a$ and $b$ to $w$ to form a $C_{3}$ on abwabw and the path cdc...d to $z$, forming an extended paw graph. In the latter case, we instead attach $a$ and $b$ to $z$ and the path on cdc...d to $w$ to again form an extended paw. Now consider broadcast initiated at the solitary leaf node. The message will reach $\gamma$ and then bounce back and forth between $\gamma$ and the $C_{3}$ for ever.
@@
Figure 8. A communication graph for Case 2 where the path cdc...d contains all remaining identifiers from $[m]$ and a non-terminating message sequence initiated at node $z$. The message sequence is given to the final unique configuration.
@@
Figure 9. Non-terminating messages sequences for cases 33 and 44 in lemma B.7. Both behave identically for the first 9 iterations, however case 33 takes two further iterations to repeat a step (indicated by dashed lines). Note that path cdc...d has been omitted here but would be attached to $z$.
Thus as we have obtained non-termination for all cases other than case 0, we must have for all $\alpha,\beta,\gamma\in[m+1]$ that $f(\alpha,\{\beta,\gamma\},\{\beta\})=\{\gamma\}$. Furthermore, we must have that $b(\alpha,\{\beta,\gamma\})=\{\beta,\gamma\}$ as otherwise the protocol would not terminate on a cycle. This also implies that $f(\alpha,\{\beta,\gamma\},\{\beta,\gamma\})=\emptyset$ to ensure termination on odd cycles when broadcast is initiated at $\alpha$. Finally, we must have that $f(\alpha,\{\beta\},\{\beta\})=\emptyset$ as $\alpha$ could be on the end of an extended paw. Thus, we have that $P$ is AF on $[m+1]$ up to degree $2$.
Fortunately, we can reuse much of the previous argument for obtaining degree $3$ behaviour. Excluding the diamond graph and the $K_{3,2}$ graph, we can use all of the constructions in Lemma B.5 by simply adding a path cdc...d to a node of degree at most $2$ which is not adjacent to $v$. However, we can slightly alter the diamond and $K_{3,2}$ graph to permit the addition of such path to a node of degree at most $2$ which is not adjacent to $v$ no matter where $v$ appears in the labelling. These are given explicitly in figure Fig. 10. Note that while, these do require $8$ nodes to construct, we can use the original constructions from Lemma B.5 until we have more than $\kappa$ nodes.
@@
Figure 11. Two graphs used in the proof of theorem 4.3 to determine identifiers’ response to receiving only a single message. In both cases cdc...d is a path containing all unused identifiers from $[m]$ where $m$ is the highest id used in labelling the body of the graph. Left: A tree that forbids sending to too small a subset of neighbours. Right: A graph that forbids sending a message to all neighbours.
@@
Figure 12. A graph used in the proof of theorem 4.3. The graph consists of a star centred at $u$ with its leaves partitioned into two sets of size $r$ and krk-r. The leaves in the set of size $r$ are connected by a binary tree of depth $\lceil\log_{2}{r}\rceil$ with root $w^{*}$ to a cycle, which in turn is connected to a path. Note that the path cdc...d contains all identifiers from $[m]$ not used in the labelling of the rest of the graph, where $m$ is the largest idid present.
@@
Figure 13. A graph used in the proof of Theorem 4.3. The graph consists of a star centred at $u$ with its leaves partitioned into three sets of size $r$, qr+1q-r+1 and kq1k-q-1. The leaves in the first two sets are then each joined to a single node labelled by ww* and vv* respectively by binary trees of depth $\lceil\log_{2}{q}\rceil$. The single nodes are connected to a node labelled $a$ which is the start of a path acdac...d that contains all identifiers in $[m]$ not used in the labelling of the body of the graph where $m$ is the largest id used in that labelling.
@@ Appendix C Uniqueness: Existence of Protocols in the relaxed cases
In this appendix, we present protocols for each of the relaxations discussed in Section 4.2
@@ C.2. Protocol with non-determinism
Random Flooding is the protocol defined by the following rule: on each round, if a node has a random bit equal to 0 it implements Amnesiac Flooding. Otherwise upon receiving a message it sends to all neighbours.
@@
Consider the random variable $X(t)$ which for $t\in\mathbb{N}$ takes the value $1$ if on rounds t,t+1,,t+Dt,t+1,...,t+D all nodes receive $1$ as a random bit, but at round $t+D+1$ all nodes receive 0, and 0 otherwise.
@@ C.3. Protocols that can read messages/send multiple messages
1-Bit Flooding is the protocol defined according to the following rule: When broadcast begins, the initiator picks 0 for the shared bit if it is a leaf and $1$ otherwise. Upon receiving a message, if the shared bit is a $1$ the node implements parrot flooding, and implements Amnesiac Flooding otherwise.
@@
If we have readable header information, then the single bit is placed in the header. On the other hand, if the nodes are able to send multiple messages per edge per round, then the initiator may choose to send $2$ messages in the case where it wishes to encode a 11 and a single one otherwise. In this case, if nodes always forward the same number of messages then the single bit is propagated and maintained.
@@
Let $S$ be a configuration of messages on $G=(V,E)$. Since messages are only reintroduced at leaves rather than on cycles or FECs, $P(S)$ is balanced on $G$ if and only if $S$ is balanced on $G$. Formally, this follows from Theorem D.5.
However, if parrot flooding is initiated at the terminal vertex of a path this configuration is balanced but will not terminate. Therefore, we require a stronger condition than balance to capture termination under parrot flooding. Consider a message path of parrot flooding (the definition of which is obtained by simply replacing the operator $A$ with $P$ in definition A.3). We immediately inherit Lemmas A.4, A.5, A.6, A.6 and A.8 as none of them depend on the behaviour of leaves. Thus any recurrent message $(u,v)\in S$ must have a message path uvwlwuvuv...wlw...uv where $l$ is a leaf node or it would be captured by the argument used in the proof of Theorem 5.1. Therefore, in fact there exists a configuration $\hat{S}=P^{k}(S)$ for some $k\in\mathbb{N}$ such that $(w,l)\in\hat{S}$ where $(w,l)$ is recurrent on $\hat{S}$. The rest of the argument is similar to that used in the proof of Theorem 5.1, and we will construct a suitable notion of balance. First, however, we must define the additional structure that produces non-termination.
@@
We have already established that balance is preserved by parrot flooding, so it remains to show that $P$-Balance is conserved on all leaf paths. Let $Q$ be a leaf path starting at $u$ and ending at $v$ on $G=(V,E)$ and let $S$ be a configuration on $G$. Then $P_{G}(S)_{Q_{p}}=P_{Q_{p},I}(S)$ for some $I\subset V$ such that $u,v\not\in I$, i.e. a step of parrot flooding on $G$ is equivalent to a step of parrot flooding on the leaf path representation with some possible external node activation. After a step of parrot flooding, two messages on $Q_{p}$ will either have the distance between their heads: stay the same, increase by two or decrease by two. Upon externally activating a non-leaf node two messages are created, with their heads an even distance apart on $Q_{p}$. Symmetrically, a collision between two messages can only be between two messages with an even distance (explicitly 0) between their heads and removes both. Thus, $P$-Balance is maintained. ∎
@@ Appendix D Proof of the Fault Sensitivity results
In this appendix, we prove the fault sensitivity and extended dichotomy results of section 5, using Theorem 5.1 as our main technical tool.
D.1. Extended Dichotomy
Here we present a proof of lemma 5.8.
Lemma D.1 (Lemma 5.8 restated).
Let $G=(V,E)$ be a graph, $S$ a configuration of messages on $G$ and $T=\{u\in V|\forall v\in N(v):(v,u)\in S\}$ the set of sink vertices. Then $A_{T,G}(\overline{A_{\emptyset,G}(\bar{S})})=S$
Proof of lemma 5.8.
Assume the message $(u,v)\in A_{T}(\overline{A(\bar{S})})$ but $(u,v)\notin S$. There are two cases. The first case is that $u$ is an initiator and therefore in $T$, however by definition the nodes in $T$ are sinks in $\bar{S}$ and so this is impossible. The second is that there exists some message $(w,u)\in\overline{A(\bar{S})}$ and $(v,u)\notin\overline{A(\bar{S})}$. For $(w,u)$ to be in $\overline{A(\bar{S})}$ requires that $(u,w)\in A(\bar{S})$, this requires some $(x,u)$ to be in $\bar{S}$. However, since $(v,u)\notin\bar{S}$ we must have that $(u,v)\in A(\bar{S})$ and so $(v,u)$ must be in $\overline{A(\bar{S})}$. Thus, we have a contradiction and so $A_{T,G}(\overline{A_{\emptyset,G}(\bar{S})})\subseteq S$
Consider a message $(u,v)\in S$ and thus, $(v,u)\in\bar{S}$. If $v$ is not a sink in $\bar{S}$, either there exists a message leaving $u$ over an edge or an edge incident to $u$ over which no message is sent in either $S$ or $\bar{S}$. In either case, $A(\bar{S})$ will include a message sent over that edge which we denote $(u,w)$. Furthermore, we know that $(v,u)\not\in\overline{A(\bar{S})}$ as that would require $(u,v)\in A(\bar{S})$ which is impossible as $(v,u)\in\bar{S}$. Since $\overline{A(\bar{S})}$ must contain the message $(w,u)$, $(u,v)$ must belong to $A(\overline{A(\bar{S})})$.
On the other hand, assume that $v$ is a sink in $\bar{S}$, then it is in $T$ and will be an initiator in $A_{T,G}(\overline{A_{\emptyset,G}(\bar{S})})=S$ unless it receives a message in $\overline{A_{\emptyset,G}(\bar{S})}$. For $v$ to receive a message in $\overline{A_{\emptyset,G}(\bar{S})}$ requires that $v$ sends a message in $A_{\emptyset,G}(\bar{S})$. However, $v$ is a sink in $\bar{S}$ and so cannot send any messages in $A_{\emptyset,G}(\bar{S})$. Thus we have a contradiction. Therefore, $S\subseteq A_{T,G}(\overline{A_{\emptyset,G}(\bar{S})})$. ∎
D.2. Fault Sensitivity
We begin with the single message failure case, restated below.
Theorem D.2 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate if either:
  • uvuv is not a bridge
  • uvuv lies on a path between odd cycles
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, uvuv is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.